# Revisiting Point Cloud Completion: Are We Ready For The Real-World?

### v2

@@ Abstract
Point clouds acquired in constrained and challenging real-world settings are incomplete, non-uniformly sparse, or both. These obstacles present acute challenges for a vital task - point cloud completion. Using tools from Algebraic Topology and Persistent Homology ($\mathcal{PH}$), we demonstrate that current benchmark synthetic point clouds lack rich topological features that are important constituents of point clouds captured in realistic settings. To facilitate research in this direction, we contribute the first real-world industrial point cloud dataset for point cloud completion, RealPC - a diverse set of rich and varied point clouds, consisting of $\sim$ 40,000 pairs across 21 categories of industrial structures in railway establishments. Our benchmark results on several strong baselines reveal a striking observation - the existing methods are tailored for synthetic datasets and fail miserably in real-world settings. Building on our observation that RealPC consists of several 0 and 1-dimensional $\mathcal{PH}$-based topological features, we demonstrate the potential of integrating Homology-based topological priors into existing works. More specifically, we present how 0-dimensional $\mathcal{PH}$ priors, which extract the global topology of a complete shape in the form of a 3-D skeleton, can assist a model in generating topologically-consistent complete shapes.
Figure 1: Qualitative comparison of synthetic and real-world object point clouds. (a) We show a complete absence of non-uniformity and noise for the synthetic PCN dataset, in contrast to multiple-levels of non-uniformity (as shown by the magnified regions) along with noise (bottom two magnifications) in the case of our RealPC dataset. (b) Comparison of Persistence Diagrams of a point cloud from RealPC vs PCN dataset. Points further away from the diagonal indicate strong topological features. RealPC has numerous significant 0 and 1-dimensional ($H_{0}$,$H_{1}$) topological features.
@@ 1 Introduction
In its simplest form, a point cloud is a discrete surface-sampling of an object or an environment, which can be further processed to generate corresponding 3D models. With the growing interest in generating such digital twins of our real-world over the years, point clouds can now be obtained using various sensors and techniques, such as LiDAR, depth cameras, photogrammetry, structured light scanning, etc. Their ability to represent complex shapes with attention to local intricate details makes them indispensable across numerous fields, such as robotics [26], autonomous driving [58], medical image analysis [62], augmented and virtual reality [35], remote sensing and geoinformatics [28], etc.
@@
Inherently, real-world point clouds are non-uniformly sparse and have rich topoloigcal features, unlike synthetic point clouds (Figure 1). They may have missing parts due to occlusions, view angle constraints, light reflectivity and object surface properties, multiple resolutions when more than one type of sensor is used, environmental factors such as fog and dust, etc. This hinders the performance of all downstream point cloud tasks, for instance, registration, surface reconstruction, object recognition, segmentation, etc. As a result, obtaining the full 3D shape representation from an incomplete and sparse point cloud plays a vital role in the practical applications of point cloud datasets. With the rise of data-driven methods, this area has been widely investigated under the umbrella term point cloud completion [21, 64, 45].
@@
Although historically motivated from a real-world perspective, as discussed above, most of the existing models in the field of machine learning-based point cloud completion report their performances on mostly synthetic datasets [48, 61, 13, 31, 52, 63]. This stems from the easy accessibility, simpler shapes, minimal to no noise, and uniformly-distributed points typically found in synthetically-generated point cloud datasets such as PCN [59], Completion3D [43], ShapeNet55/34 [56], MVP [36]. However, due to the excessive tailoring of these models to such synthetic datasets, they fall short of providing similar results for realistic datasets, as we will see in the upcoming sections.
@@
The unordered representation of a point cloud does not allow for the encoding of any structural information. This makes point cloud processing very challenging. Persistent Homology has been show to learn global structural properties of datasets. It is a powerful tool from topological data analysis (TDA) that has gained traction in 3D vision applications, particularly for analyzing and understanding the underlying shape and structure of data from different modalities - point clouds, images, graphs, [42, 34, 8] etc. It helps in identifying topological features, such as connected components, loops, and voids, across multiple scales. However, as far as our knowledge goes this line of research has not yet been explored for the task of completing point clouds in general, let alone real-world ones.
@@ 2.1 Benchmark Object-Level Point Cloud Datasets
Not only is the processing of real-world point cloud data arduous but also its acquisition comes with its own set of challenges. With the availability of high-end 3D-scanning devices nowadays, this task ultimately boils down to balancing the time involved, the accuracy required, and the setup costs. In addition, there is no particular standard storage format followed by the available sensors in the market today. Moreover, such point cloud datasets have background outliers which can make most of the modern data-driven algorithms blow up and therefore demand manual outlier removal, which is extremely time-consuming. All of these factors combined pose a great difficulty in accessing and working with realistic point cloud data and introduce an inevitable hurdle for 3D vision research community. Hence, researchers prefer working on synthetically generated datasets instead. We now provide a brief discussion on all existing benchmark datasets on which recent SoTA point cloud completion models have experimented on.
The PCN dataset [59] contains paired partial and complete point clouds derived from 30,974 CAD models across 8 categories from the ShapeNet repository [11]. Two more datasets derived from the same CAD repository [11] are ShapeNet55 and ShapeNet34 datasets [56]. ShapeNet-55 includes 52,470 objects from 55 categories. The ShapeNet-34 dataset is derived by splitting the original ShapeNet-55 dataset into 34 seen categories and 21 unseen categories. Another synthetic dataset is the MVP dataset [36], which consists of more than 100,000 point clouds from 16 categories sampled again from CAD models.
Although fewer in number, some recent real-world point cloud object datasets include OmniObject3D [51], MatterPort3D [10], ScanNet [17], and KITTI [25]. The ScanNet and MatterPort3D datasets offer 3D reconstructions of indoor spaces, while the KITTI dataset captures scenes from outdoor environments. All of these datasets are unpaired and hence contain only partial shapes and no ground truth, naturally making them infeasible for supervised training.
@@ 2.2 Point Cloud Completion
Point cloud completion has been extensively studied on synthetic point clouds (e.g. PCN, Shapenet, etc). Most works approach this it as a point features learning and per point generation problem. Voxel-based and point-based approaches have been primarily used for this task. GRNet [54] utilizes 3D CNNs and multilayer perceptrons (MLPs) for point cloud representation. However, voxel-based methods struggle with limited input points, resulting in incomplete voxel neighborhoods and hole-ridden outputs. This robustness limitation hinders their effectiveness. Point-based methods directly operate on the raw points. Started with PCN [59], several works [52, 50, 63] have used this strategy for effective completion. ODGNet [6] utilizes a Seed Generation U-Net for enhanced seed point representation. It incorporates Orthogonal Dictionaries to learn shape priors, compensating for missing information during inference. PointAttN [49] leveraging cross-attention and self-attention mechanisms, incorporating Geometric Details Perception (GDP) and Self-Feature Augment (SFA) blocks to establish structural relationships among points. PointTr [57] represents point clouds as unordered groups with position embeddings, converting them to point proxy sequences for generation via a transformer encoder-decoder architecture. There have been several such works [41, 60, 44, 37, 53, 50, 46, 30, 3, 55, 2] most of which revolve around the point feature learning-based encoder-decoder architecture. The current literature works well on synthetic point clouds, but their performance on real-world point clouds needs to be probed. We follow this line of work and demonstrate some interesting and contrasting observations in this direction.
@@ 2.3 Persistent Homology and Topological Deep Learning
Persistent Homology ($\mathcal{PH}$) has its origins in algebraic topology and enables the measurement of topological features which characterize the shape of the data. Persistent Homology can be seen as an adaption of homology  [20, 38, 5] in case of point clouds. It captures the evolutions of topological features (cycles, voids, etc.) in several dimensions (0, 1, 2) in a dataset. $\mathcal{PH}$ is concreted with the study of invariants across multiple scales. Characterizing these features allows $\mathcal{PH}$ to discern essential structures and patterns that is not possible using conventional methods. Topological methods have been recently used across several learning-based tasks. Several works [27, 39, 29, 14, 32, 12, 18] have used topological tools to augment learning with better feature descriptors. Vectorized topological descriptors extracted using $\mathcal{PH}$ have been used for interpretability of models [22], adversarial learning. [24] introduce topological representation for auto-encoding and demonstrate the topological richness of the learned representations. $\mathcal{PH}$ has also been used for several geometric problems like - surface reconstruction, pose matching [9, 4, 18] etc. Several works have also proposed general-purpose differentiable topology layers to extract descriptors for topological regularization [23].
@@ 3 Background
Figure 2: Left to Right (a to c). Progression of filtration on a point cloud over different spatial resolutions as the distance threshold increases [33]. Creation and destruction of $k-dim$ topological features documented in (d) which is a persistence diagram representing $(b_{i},d_{i})$ pairs - each point corresponds to a $k-dim$ homology that takes birth at $b_{i}$ and dies at $d_{i}$.
A $k$-simplex is the convex hull of $k+1$ affinely independent points, forming a line ( in 1D), triangle (in 2D), or a tetrahedron (in 3D), etc. A simplicial complex is a collection of simplices with simplex intersections. In topology, homology counts $k$-dimensional holes(0-D: connected components, 1-D: loops 2-D: voids), etc.
@@
4 Object-Level Training Dataset Creation
@@
In light of the challenges outlined in Section 2.1 concerning the acquisition of real-world point cloud datasets, in this section we demonstrate a methodology (Figure 4) to extract and process paired object point cloud datasets from existing open-source scene-level datasets. Their paired nature ensures suitability for any supervised point cloud completion algorithm, therefore, lowering the threshold for experimenting on real-world data. Although we have explored this methodology in a railway-environment setup, this exact procedure can be followed for any scene-level point cloud dataset.
We use four open-source railway datasets: [15] by Hungarian State Railways acquired with a Riegl VMX-450 high-density mobile mapping system; [40] by Wuhan University, in which urban railway dataset was captured using Optech’s Lynx Mobile Mapper System, rural railway dataset with MLS system equipped with HiScan-Z LiDAR sensors, and plateau railway dataset with Rail Mobile Measurement System (rMMS) equipped with a 32-line LiDAR sensor; [1] by SNCF Réseau, the French state-owned railway company; and a Catenary arch dataset [47] given by Strukton Rail, captured using Trimble TX8 Terrestrial Laser Scanner (TLS). A few scenes from these datasets have been visualized in Figure 3.
The mentioned scene-level point clouds are pre-annotated into relevant sections such as vegetation, overhead cables, railway tracks, industrial support and transmission structures, ground, etc. For this work, we extract industrial structures from these scene-level point clouds. As a result, we obtain a point cloud with multiple industrial structures as shown in Figure 4 Input to (A). Referring to the same figure, our methodology is divided into five main parts. In (A) we employ HDBSCAN [7], a hierarchical clustering algorithm, to cluster individual industrial structures. HDBSCAN works accurately for noisy and complex data and automatically detects the number of data clusters. In (B), a manual inspection is performed to extract industrial structures which can serve as good ground truth for supervised training. Finally, in (C), (D), and (E), we process these ground truth point clouds in three different ways to induce uniform sparsity, non-uniform sparsity, and incompleteness as explained in detail ahead.
For making a partial point cloud (C), we pick a random viewpoint around a ground truth point cloud and remove N number of points that are farthest away from this viewpoint. Similarly for sparsifying the same point cloud non-uniformly (D), we again pick a random viewpoint but this time we assign a probability to each point, which is either proportional or inversely proportional to a point’s cubed distance to this viewpoint. Then, we sample N points based on these probability values. This is repeated for different values of N and for different viewpoints. For uniform sparsification (E), we randomly sample N points and then repeat for different values of N. Simultaneous to these steps, we perform manual noise-removal. These three steps are conducted for all ground truth point clouds to obtain a paired object-level training dataset.
As a final step, we manually examine all the industrial structures and classify them into 21 classes based on their geometric shapes. From now on we will call this dataset RealPC. On publication, to help advancement in this line of research, we will make this processed object-level training dataset as well as the code for this processing pipeline open-source.
Figure 3: Scene-level point clouds from different acquisition techniques (a) [1], (b) [15], (c) [40], and (d) [47]
@@
Figure 4: Object-level training dataset creation methodology. Input to (A) are the segmented industrial structures from scene-level point clouds. Figure (B) shows how manual inspection was done to extract ground truth data. Output of (C), (D), and (E) are the 3 variants of sparse and incomplete point clouds
@@
Class 1 Class 2 Class 3 Class 4 Class 1 Class 2 Class 3 Class 4 Class 1 Class 2 Class 3 Class 4
PCN 10.6 8.8 8.6 14.0 18.1 12.0 14.7 20.4 79.5; 34.2 52.4; 22.2 66.0; 31.9 87.1; 36.9
ShapeNet 9.8 12.2 14.9 11.2 13.9 14.6 20.2 17.8 44.4; 18.6 47.2; 20.2 66.0; 29.4 56.4; 26.5
RealPC 137.4 99.2 93.1 92.5 234.2 222.4 128.1 148.2 382.3; 155.7 322.5; 114.9 341.2; 149.5 274.8; 126.4
Table 1: Quantitative comparative analysis of RealPC against PCN and ShapeNet using three metrics - sum of 0 and 1-dimensional persistence of topological features, non-uniformity, and noise. Best values are bold-faced. All values are in units of $10^{-4}$. All numbers demonstrate that RealPC has stronger Homology-based topological features, non-uniformity and noise. For a visual demo see Fig. 1

### v3

@@ Abstract
Point clouds acquired in constrained, challenging, uncontrolled, and multi-sensor real-world settings are noisy, incomplete, and non-uniformly sparse. This presents acute challenges for the vital task of point cloud completion. Using tools from Algebraic Topology and Persistent Homology ($\mathcal{PH}$), we demonstrate that current benchmark object point clouds lack rich topological features that are integral part of point clouds captured in realistic environments. To facilitate research in this direction, we contribute the first real-world industrial dataset for point cloud completion, RealPC - a diverse, rich and varied set of point clouds. It consists of $\sim$ 40,000 pairs across 21 categories of industrial structures in railway establishments. Benchmark results on several strong baselines reveal that existing methods fail in real-world scenarios. We discover a striking observation - unlike current datasets, RealPC consists of multiple 0- and 1-dimensional $\mathcal{PH}$-based topological features. We prove that integrating these topological priors into existing works helps improve completion. We present how 0-dimensional $\mathcal{PH}$ priors extract the global topology of a complete shape in the form of a 3D skeleton and assist a model in generating topologically consistent complete shapes. Since computing Homology is expensive, we present a simple, yet effective Homology Sampler guided network, BOSHNet that bypasses the Homology computation by sampling proxy backbones akin to 0-dim $\mathcal{PH}$. These backbones provide similar benefits of 0-dim $\mathcal{PH}$ right from the start of the training, unlike similar methods where accurate backbones are obtained only during later phases of the training.
Figure 1: Qualitative comparison of current existing and RealPC object point clouds. (a) We show a complete absence of non-uniformity and noise for the PCN dataset [64], in contrast to multiple levels of non-uniformity (as shown by the magnified regions) along with noise (bottom two magnifications) in the case of our RealPC dataset. (b) Comparison of Persistence Diagrams of a point cloud each from RealPC and PCN dataset. Points further away from the diagonal indicate strong topological features. Unlike PCN, RealPC has numerous significant 0- and 1-dimensional ($H_{0}$, $H_{1}$) topological features.
@@ 1 Introduction
In its simplest form, a point cloud (PC) is a discrete surface-sampling of an object or an environment. With the growing interest in digital twins of the real-world, PCs can now be obtained using various sensors and techniques, such as LiDAR, depth cameras, photogrammetry, structured light scanning, etc. Their ability to represent complex shapes with attention to local intricate details makes them indispensable across numerous fields, such as robotics [26], autonomous driving [63], medical image analysis [67], augmented and virtual reality [38], remote sensing and geoinformatics [28], etc.
@@
Inherently, real-world PCs are non-uniformly sparse, noisy, complex-structured and topologically rich, unlike their synthetic counterparts (Figure 1). Moreover, they may also have missing parts due to occlusions, view-angle constraints, object surface properties, environmental factors, etc. This hinders the performance of all downstream PC tasks, for instance, registration, surface reconstruction, object recognition, segmentation, etc. As a result, obtaining the full 3D shape representation of an incomplete and sparse PC plays a vital role in the practical applications of PC datasets. With the rise of data-driven methods, this area has been widely investigated under the umbrella term point cloud completion [21, 69, 47].
@@
Although historically motivated from a real-world perspective, as discussed above, most of the existing models in the field of machine learning-based PC completion report their performances on synthetic datasets [51, 66, 13, 33, 56, 68]. This stems from the easy accessibility, simpler shapes, minimal to no noise, and uniform point distribution of controlled and synthetically-generated PC datasets [64, 45, 62, 39]. However, due to the excessive tailoring of these models to these existing datasets, they fall short of providing similar results for real-world datasets.
@@
The unordered representation of a PC itself does not allow for the encoding of any structural information, which makes PC processing very challenging. Persistent Homology ($\mathcal{PH}$), a powerful tool from Topological Data Analysis (TDA), has proven to be highly effective in extracting structural properties from 3D datasets. It has gained traction in vision applications, particularly for analyzing and understanding the underlying shape and structure of data from different modalities such as PCs, images, graphs [44, 37, 8], etc. However, as far as our knowledge goes this line of research has not yet been explored for the task of completing PCs in general, let alone real-world ones. Hence, this paper presents the following key contributions:
  • We present a real-world paired (ground truth complete and sparse/partial incomplete) industrial object PC dataset RealPC, captured in challenging, uncontrolled and realistic multi-sensor settings.
  • We demonstrate that RealPC PCs are intrinsically different from existing synthetic ones via three important metrics, $\mathcal{PH}$ features, non-uniformity and noise. Using tools from TDA, we discover that they have rich and versatile topological features. Benchmarking exercises on non-neural and neural methods reinforce our discovery and establish that current methods for existing datasets do not work well with such realistic ones.
  • We prove the significance of integrating 0-dim $\mathcal{PH}$ priors into existing completion models (under certain conditions). Since Homology extraction is costly, we design a simple and intuitive Homology Sampler network BOSHNet, which extracts multiple Homology backbones from a PC at various scales. These priors force a network from the start of the training to stick to the backbone while generating complete shapes, enabling precise completion.
@@ 2.1 Benchmark Object-Level Point Cloud Datasets
Not only is the processing of real-world PC data arduous but also its acquisition comes with its own set of challenges. With the availability of high-end 3D-scanning devices nowadays, this task ultimately boils down to balancing the time involved, the accuracy required, and the setup costs. In addition, there is no particular standard storage format followed by the available sensors in the market today. Moreover, such PC datasets have background outliers which can make most of the modern data-driven algorithms blow up and therefore demand manual outlier removal, which is extremely time-consuming. All of these factors combined pose a great difficulty in accessing and working with realistic PC data and introduce an inevitable hurdle for the 3D vision research community. Hence, researchers prefer working on synthetically-generated datasets instead.
Among some of the well-known datasets for PC completion is the PCN dataset [64], which contains paired partial and complete PCs derived from 30,974 CAD models across 8 categories from the ShapeNet repository [11]. Two more datasets derived from the same CAD repository [11] are ShapeNet55 and ShapeNet34 datasets [62]. ShapeNet-55 includes 52,470 objects from 55 categories. The ShapeNet-34 dataset is derived by splitting the original ShapeNet-55 dataset into 34 seen categories and 21 unseen categories. Another such dataset is the MVP dataset [39], which consists of more than 100,000 PCs from 16 categories sampled again from CAD models.
Although fewer in number, some recent real-world PC object datasets include OmniObject3D [55], MatterPort3D [10], ScanNet [16], and KITTI [25]. The first three datasets offer PCs captured in extremely controlled environments, in contrast to real-world settings, which induce numerous factors of variation, disturbance and noise. Moreover, all of these datasets are unpaired and hence contain only partial shapes without ground truth, making them infeasible for supervised training.
@@ 2.2 Point Cloud Completion
PC completion has been widely explored for existing datasets such as PCN and ShapeNet. Most approaches frame this as learning point-wise features and generating points individually to reconstruct the complete structure. Voxel and point-based techniques have been used extensively for this purpose. GRNet [58] uses 3D CNNs and MLPs for PC representation. Voxel-based methods struggle with limited input points, resulting in incomplete neighborhoods and hole-ridden outputs. Point-based methods operate directly on raw PCs (e.g. PCN [64]). Multiple works [56, 54, 68] have used this strategy for effective completion. ODGNet [6] incorporates a seed generation U-Net for enhanced seed representation, and orthogonal dictionaries to learn shape priors. PointAttN [51] leverages cross and self-attention mechanisms, using Geometric Details Perception and Self-Feature Augment blocks to establish structural relationships between points. PointTr [62] represents PCs as unordered groups with position embeddings, converting them to point proxy sequences for generation via a transformer encoder-decoder architecture. There are several similar works [43, 65, 46, 39, 57, 54, 48, 30, 3, 59, 2], which revolve around the point feature learning approach. The current literature works well on these PCs, but their performance on real-world PCs remains relatively less explored.
@@ 2.3 Persistent Homology and Topological Deep Learning
Persistent Homology ($\mathcal{PH}$), having its origins in algebraic topology, measures topological features, which characterize the shape of the data. These features help in discerning essential structures and patterns, which conventional methods cannot. Topological methods have been recently used across several learning-based tasks to augment architectures with better feature descriptors [27, 41, 29, 14, 35, 12, 17]. Vectorized topological descriptors extracted using $\mathcal{PH}$ have been used for interpretability of models and adversarial learning [22]. [24] introduces topological representation for auto-encoding and demonstrates the topological richness of the learned representations. $\mathcal{PH}$ has also been used for several geometric problems like surface reconstruction, pose matching [9, 4, 17], etc. [23] proposes general-purpose differentiable topology layers to extract descriptors for topological regularization.
@@ 3 Background
Let $X$ be some general bounded surface in 3-space, discretely sampled to form a point cloud, represented by the unordered set $\phi$. A standard tool from algebraic topology represents the topology of $X$ by a simplicial complex ${\cal C}$, a finite set of simplices of different dimensions. In our case the surface $X$ has dimension 2, so the only simplices we need are vertices (the points $\phi$), edges (connecting some pairs of $\phi$) and triangles (bounded by some triples of edges). By definition, the different simplices of this complex are either disjoint or intersect in a common subsimplex. The idea is that ${\cal C}$ provides a triangulation of $X$ such that the union of all simplices is topologically equivalent (homeomorphic) to $X$. The nice thing about representing a space $X$ by a simplicial complex ${\cal C}$ is that we can consider integer linear combinations of simplices at each dimensional level, yielding chains or cycles of edges and triangles, including so-called boundary maps to obtain cycles of one dimension less. The benefit of this added algebraic structure is the homology of the complex ${\cal C}$, identifying vertices that are linked by edges, and detecting cycles of edges that bound holes of the surface, yielding topological features of $X$, independent from the chosen triangulation ${\cal C}$.
Figure 2: (a) to (c) Progression of filtration on a PC over different spatial resolutions as the distance threshold increases [36]. (d) Birth and death of $k$-dim topological features documented in the form of a persistence diagram, i.e., $(b_{i},d_{i})$ pairs, so that each point corresponds to a homology which is born at $b_{i}$ and dies at $d_{i}$.
@@
4 RealPC Dataset
4.1 Creation
@@
In light of the challenges outlined in Section 2.1 concerning the acquisition of real-world PC datasets, in this section we demonstrate a methodology (Figure 4) to extract and process paired object PC datasets from existing open-source scene-level datasets. Their paired nature ensures suitability for any supervised PC completion algorithm, therefore, lowering the threshold for experimenting on real-world data. Although we have explored this methodology in a railway-environment setup, this exact procedure can be followed for any scene-level PC dataset.
We use four open-source scene-level annotated railway datasets from different countries, a few of which are visualized in Figure 3 (more about them in supplementary material). We extract annotated industrial structures from these scene-level PCs, hence obtaining a PC with multiple industrial structure shapes (Figure 4 Input to (A)). Our creation pipeline (Figure 4) is divided into five parts. (A) We deploy HDBSCAN [7], a hierarchical clustering algorithm to cluster individual structures. (B) We perform manual inspection to extract complete industrial structures, which can serve as good ground truth (GT). Steps (C), (D), and (E) process these GT PCs in three different ways to induce uniform sparsity, non-uniform sparsity and incompleteness: (C) pick a random viewpoint around a GT PC and remove N points farthest away from this viewpoint, (D) pick a random viewpoint around a GT PC, and sample N points w.r.t the probability values assigned to the points, which are either proportional or inversely proportional to a point’s cubed distance from this viewpoint, and (E) randomly sample N points from a GT PC. While (C) and (D) are repeated for multiple viewpoints and N values, (E) is performed for several values of N. We manually examine all the industrial structures and classify them into 21 classes based on their geometric shapes, hence obtaining a real-world PC dataset which we call RealPC.
We shall make RealPC as well as the code for the processing pipeline open-source.
Figure 3: Scene-level PCs from different acquisition techniques (a) [1], (b) [15], (c) [42], and (d) [49].
@@
Figure 4: Object-level training dataset creation methodology. Input to (A) shows the segmented industrial structures from scene-level PCs. Figure (B) shows how manual inspection was done to extract ground truth data. Output of (C), (D), and (E) are the three variants of sparse and incomplete PCs.
@@
Class 1 Class 2 Class 3 Class 4 Mean Class 1 Class 2 Class 3 Class 4 Mean Class 1 Class 2 Class 3 Class 4 Mean
PCN 10.6 8.8 8.6 14.0 12.2 18.1 12.0 14.7 20.4 19.7 79.5; 34.2 52.4; 22.2 66.0; 31.9 87.1; 36.9 86.0; 37.5
ShapeNet 9.8 12.2 14.9 11.2 23.7 13.9 14.6 20.2 17.8 31.8 44.4; 18.6 47.2; 20.2 66.0; 29.4 56.4; 26.5 101.3; 41.0
RealPC 137.4 99.2 93.1 92.5 113.7 234.2 222.4 128.1 148.2 173.8 382.3; 155.7 322.5; 114.9 341.2; 149.5 274.8; 126.4 345.2; 155.5
Table 1: Quantitative comparative analysis of RealPC against PCN and ShapeNet with respect to Noise (×$\times 10^{-4}$), Non-uniformity (×$\times 10^{-4}$), and 0 and 1-dimensional persistence of topological features (×$\times 10^{-4}$). All numbers demonstrate that RealPC has stronger homology-based topological features, and higher non-uniformity and noise. Refer to Figure 1 for a qualitative comparison.